OpenStudy (anonymous):

Solve the system. 2y = 4x + 2 2y = -x + 7

OpenStudy (anonymous):

subtract eq1 with eq2 4x+2-2y - -x+7-2y -------- 3x-5=0 3x=5 x=5/3 plug into eq1 2y=4(5/3)+2 y=(4(5/3))/2+1 y=4.333

OpenStudy (anonymous):

oh forgot to mention, using matrices, sorry.

OpenStudy (anonymous):

Well... that certainly changes things.... 2y = 4x + 2 2y = -x + 7 Matrix A 2 2 Matrix B 4 2 -1 7 Inverse of A = 2 2 multiply A with B result: x=6 y=18 (This may be wrong idk)

OpenStudy (anonymous):

yea, i'm pretty sure it's wrong.